OpenStudy (anonymous):

I have an arc length of 1/2 how many degrees is that the circumference of the circle is 45

OpenStudy (aum):

The arc length is proportional to the angle it makes at the center. Let x degrees be the angle it makes at the center. x / arc length = 360 / circumference x / (1/2) = 360 / 45 2x = 8 x = 4 degrees.

jimthompson5910 (jim_thompson5910):

Another way to look at it. The arc length is (1/2)/45 = 1/90 of the entire circumference So the central angle is also 1/90 of a full revolution/rotation (1/90)*(360) = 4 degrees is the central angle